# Adversarial Attacks on Reinforcement Learning Agents for Command and Control

### v1

@@ 1 Introduction
Deep Reinforcement Learning (DRL) has been successfully used to train agents in several tactical and real-time strategy games such as StarCraft[1] and DoTA[2], which involve complex planning and decision-making. These agents have demonstrated proficiency in coming up with winning strategies comparable to that of experienced human players (AlphaStar [3], OpenAI Five[4]) through techniques like self–play, imitation learning, etc. As a result, in recent years, there has been mounting interest in the military research community in applying these RL techniques to tasks such as operational planning and command and control (C2). Simultaneously, traditional game engines have been repurposed to facilitate automated learning (pySC2[5], SMAC[6, 7], pyDoTA[8]) and new ones developed for battlefield simulation [9, 10, 11, 12, 13], creating what are effectively digital wargames. The driving force behind this research has been to improve and augment strategies used on the battlefields of the future, which are expected to be more complex and unconventional, possibly beyond the cognitive abilities of a human commander.
@@
Recent works [14] have had considerable success in winning simulated wargames using C2 agents that have been trained through reinforcement learning techniques and synthetic data. This has been possible partly due to the scalability of RL training which has proved to be a massive advantage for exploring and exploiting different strategies when faced with difficult or complicated scenarios and only partial information about the environment. However, these evaluations are done in benign environments where information available to the C2 agent is assumed to be uncorrupted. Realistically, this is unlikely in battlefield situations as information there may have inherent noise because of the mode of collection (from sensors or other input source) or may be tampered with by enemy forces. In this work, we evaluate the robustness of such a trained agent when subject to potentially adversarial inputs in the context of C2.
In order to do so, first, we use the StarCraft II Learning Environment (SCLE[15]) to model conflict between two teams, the BlueForce and the RedForce. The C2 agent directs the BlueForce to win the battle by eliminating the RedForce troops. Next, we assume the an attacker present in the environment tampers with the observations collected from the battlefield before they are made available to the C2 agent. The added perturbations termed adversarial perturbations are constructed to be highly imperceptible to evade detection while maximally subverting the C2 agent’s policy to something deleterious (Figure 1). We then evaluate the resulting drop in performance of the agent on several metrics as well as analyze the deviation in the course of action from a military perspective.
@@
Our evaluations demonstrate the susceptibility of vanilla RL training to adversarial perturbations and the need for robust training mechanisms and sophisticated detection and prevention techniques especially for such critical scenarios.
@@ 2.1 RL for Command and Control (C2)
Mission success in military C2 requires disseminating control actions to designated forces based on real-time Intelligence, Surveillance, and Reconnaissance (ISR) information of the operational environment, such as kinetic and non-kinetic (e.g., weather, political, economic, cultural) variables, and terrain information. Before the battle begins, the commander and their specialized staffing officers must develop a detailed mission plan encapsulated as a Course Of Action (COA). The development of a COA requires a detailed analysis of the operational environment, predictions of the opposing force’s COA (strategy), and wargaming to identify a friendly force COA that is finely tuned to meet the mission requirements [16]. Typical military planning is solely based on the commander and staffing officers, and their allocated time before the battle dictates the number of possible COAs that can be considered. Additionally, each potential COA must be wargamed and fine-tuned against a small set of opposing force COAs to identify strengths and weakness. This can result in suboptimal (heuristic-based) COA [17].
To circumvent this limitation, future military planning is envisioned to incorporate an Artificial Intelligent (AI) commander’s assistant that can generate and recommend COAs to aid in the military planning process. Recent developments in deep RL for strategy games provides a promising direction to develop control policies for C2 [3]. The algorithms developed allow an AI agent to learn the best control policy that optimizes a predefined reward function by playing millions of simulated games through the exploration of many environmental state and action pairs. To extend these formulations for military C2 requires modeling, simulating, and wargaming a large number of battles faster than real-time in a virtual environment that emulates realistic combat characteristics. Furthermore, the existing RL algorithms must be adapted to handle a large number of heterogeneous actors, doctrine based control strategies, complex state and action space, and uncertain information.
Previous work on RL for C2 found that the StarCraft II gaming engine, developed by Blizzard Entertainment [1], provides a simulation environment that can be militarized and used for prototyping an AI commander’s assistant [18, 14]. The following subsections provide details of the StarCraft II environment and the scenarios considered in this paper.
@@ 2.2 StarCraft II C2 Environment
StarCraft II is a multi-agent real-time strategy game developed by Blizzard Entertainment [1] that consists of multiple players competing for influence and resources with the ultimate goal of defeating the other players. As a collaboration between Deepmind and Blizzard Entertainment, the StarCraft II Learning Environment (SC2LE) Machine Learning API was developed to allow researcher to study many difficult challenges associated with RL [5]. For example, controlling heterogeneous assets in a complex state and action space with uncertainties. Furthermore, the StarCraft II Editor allows developers to construct custom scenarios, making it possible to develop RL agents for C2.
Previous work extended this framework to militarize the SC2LE and develop a C2 simulation and experimentation capability that interfaces with deep RL algorithms via RLlib [18], an open-source industry-standard RL library [19]. The icons were re-skinned to portray standard military symbology, new StarCraft II maps were designed to emulate realistic combat characteristics, such as environment terrain and asset attributes (e.g., visibility, weapons, weapons ranges, and damage), and a custom wrapper was designed to train RL agents. For a detailed description of this framework, see [18]. In our work, we use this framework to develop baseline RL agents for C2 and study the effects of adversarial attacks on the learned policies in the following two scenarios.
2.2.1 Custom StarCraft II Scenarios
@@
TigerClaw: The TigerClaw melee map (Figure 3) is a high-level recreation of the TigerClaw combat scenario (Figure 2) developed using the StarCraft II map editor. The scenario was developed by Army subject-matter experts (SMEs) at the Captain’s Career Course, Fort Moore, Georgia. The BlueForce is an Armored Task Force (TF) which consists of combat armor with M1A2 Abrams, mechanized infantry with Bradley Fighting Vehicles (BFV), mortar, armored recon cavalry with BFV, and combat aviation. The RedForce is a Battalion Tactical Group (BTG) with attached artillery battery and consists of mechanized infantry with BMP, mobile artillery, armored recon cavalry, combat aviation, anti-armor with anti-tank guided missiles (ATGM), and combat infantry.
@@
National Training Center (NTC): The NTC map is a representation of the Ft. Irwin and Bicycle Lake area as seen in Figure 4. The Blue and RedForce units are the same as TigerClaw, but there is also a capability to randomize the starting forces. RedForce initially begins in the Bicycle Lake region, while BlueForce is set up in defense of Fort Irwin. Thus the RedForce will maneuver from the Bicycle Lake area and attempt to destroy the BlueForce in the Ft. Irwin area. The goal of this map was to investigate the impact of new terrain and to determine if the BlueForce would adopt a defensive strategy.
@@
2.2.2 Reinforcement Learning Algorithms
Asynchronous Advantage Actor-Critic (A3C) [20]: A3C is an asynchronous version of the Advantage Actor-Critic (A2C) algorithm for RL that uses multiple agents to learn a policy and an estimate of the value function in parallel. At each timestep, the agent takes an action based on the current policy and receives a reward while transitioning to a new state. It then calculates the advantage function and updates its local copies of the actor-critic networks. This process takes place independently and asynchronously for each agent. A central parameter server that stores the global parameters of the networks is updated periodically by each agent and is used to initialize the local parameters of the agents’ local networks. The asynchronous nature of the updates leads to more efficient exploration of the environment and reduces the correlation between the updates leading to more stable and efficient learning. Typically, the policy and value functions are parameterized by a shared neural network with a softmax output for the policy and a linear output for the value function.
Proximal Policy Optimization (PPO) [21]: PPO is a policy gradient method that aims to improve the performance and stability of trust-region[22] methods by introducing a clipped surrogate objective function. This objective function effectively restricts policy change to a small range thus reducing the variability in training of the actor. During training, the PPO algorithm iteratively updates the actor as well as the action value function and state value function using the temporal-difference (TD) method. Further, this objective function enables PPO to guarantee monotonic improvements in the objective. This allows for faster convergence without strict constraints leading to more accurate and stable performance of the agent.
@@ 3.1 State and Action Space
The action space within the StarCraft II environment is large since it is a combination of the following three components: the number of units available to the commander, the number of possible actions that each unit can execute, and the $(x,y)$ pixel location within the size $256\times 256$ minimap where the action will be executed. To reduce the action space, the StarCraft II C2 environment first restricts the number of units by defining control groups, which lumps common units together to reduce the overall number of units needed to be controlled. In our custom scenarios, the control groups are defined as – 1 BlueForce: “AVIATION", “MECH_INF", “MORTAR", “SCOUT" and “TANK" 2 RedForce: “ANTI_ARMOR", “ARTILLERY", “AVIATION", “INFANTRY" and “MECH_INF"
Within each control group, we restrict the possible actions to be either “NO_OP" or “ATTACK(x,y)", where the function of the former is to essentially do nothing, while the latter moves the control group (i.e., all associated units) to a desired $(x,y)$ location and attacks any enemy within its firing range along the way. Additionally, the number of $(x,y)$ pixel locations is reduced by segmenting the minimap into nine disjoint quadrants with locations defined as “LEFT", “CENTER", and “RIGHT" for the $x$-axis, and “TOP", “CENTER", and “BOTTOM" for the $y$-axis, where the exact pixel location is the center of the quadrant.
3.1.1 Reward Structure
TigerClaw: The reward function for the TigerClaw map consists of -
@@ 3.1.1 Reward Structure
NTC: The NTC map shares the same attrition reward function as TigerClaw, but the terrain rewards are not included. That is,
@@ 3.2.1 The Policy Network
The C2 agent in this paper uses the same policy network as our previous work [14]. The network takes as inputs, three kinds of observations and the control group
@@ 4 Adversarial Attacks on RL agents
Prior works on the robustness of RL training have focused on evaluating the algorithms from a perspective of sensitivity to environment dynamics [23, 24] or the ability to train adaptive adversarial policies against them [25, 26]. In this work, we focus on the former approach. A lot of previous research has shown that neural network predictions are highly sensitive to perturbations in their input space[27, 28, 29]. As DRL approaches typically rely on parameterizing policies with neural networks, they suffer from the same vulnerabilities.
@@ 4.1 Adversarial Attacks on Image Classifiers
4.1.1 Fast Gradient Sign Method (FGSM) [28]
@@
where $y_{t}$ is the ground truth label of the sample $x$ and $\varepsilon$ is the perturbation budget that controls the amount of distortion to the original sample. In the case of FGSM, it bounds the $l_{\infty}$ norm of the perturbation added to the original image. Intuitively, the attack moves the sample $x$ in the direction of $\nabla_{x}L(f(w;x),y_{t})$ which maximizes the classifier loss $L$. In this work, we prefer FGSM to more powerful attacks such as PGD[29] and C&W[27] because of its lower computational cost, which leads to a more efficient attack.
@@ 4.2.1 Threat Model
4.2.2 Attack Methodology
We use FGSM as a basis for generating perturbations at inference time and use the modification presented in [23] to target policies instead of classifiers. Unlike supervised learning, we do not know the “ground truth" action at any given timestep in RL and we assume that the action predicted by the policy network with the highest likelihood is optimal.
@@ 5.1 Experimental Setup
Our experiments were conducted on a cluster node with two AMD Epyc 7763 “Milan" CPUs 128 cores with 256GBs of main memory. We used the SC2 framework described in Section 2.2 to train our RL agents. Each training run was performed using 90 workers, each occupying a single core (90 parallel instances of the game) and consumed about 170GBs of memory. The PPO and A3C agents were trained for 5K/25K iterations respectively corresponding to around 40M/100M timesteps or 134K/376K episodes respectively. The rollouts were visualized using a custom pygame interface.
@@ 5.2 Evaluation of Inference Time Attacks on Agent Reward
5.2.1 Vulnerability to Adversarial Perturbations
@@
5.2.2 Analyzing Agent Behavior under Attack
@@
To better understand this shift in behavior, we compare actions taken by the agent in a benign environment to those taken by the same agent when an adversary is present. That is, at each timestep we compute the actions that would be taken by the agent if the observations were maliciously modified and compare these “subverted actions" to the actions that are actually taken by the agent at that timestep. We do so by perturbing the observations using our attack (with $\varepsilon=0.1$) and using the agent to predict the next action. The normalized action frequencies are plotted in Figure 10 for a PPO/TigerClaw agent over 5 episodes. A clear difference is observed in the actions taken in the two cases. Actions $(0,2,1),(1,1,1),(1,2,1)$ are sampled frequently when perturbed observations are presented to the agent. These actions misdirect the BlueForce as they target areas on the map that do not contain any RedForce troops.
@@
5.2.3 Strength & Reliability of the Attack
@@
Attacking the A3C/TigerClaw agent: As can be seen in Figure 7(b), the attack appears to fail on the C2 agent trained using A3C on the TigerClaw map. Even for large perturbations ($\varepsilon>0.1$), the attack fails to degrade the agent’s reward as extensively as the other cases. Interestingly, we note that even in a benign environment, the A3C/TigerClaw agent achieves a significantly lower reward than the PPO/TigerClaw agent.
@@
5.2.4 Component-wise Impact of Input on the Attack
@@
Figures 16(a), 16(b), and 16(d) indicate that for smaller perturbations levels ($\epsilon<0.5$), modifications to the screen component are primarily responsible for a successful attack. Further, we only see a minor degradation in the reward for small nonspatial-only perturbations in both scenarios. On the other hand, the drop in mean reward because of screen-only perturbations is similar to that in the baseline attack. Intuitively, this skewed importance makes sense as the screen component is a much larger input component ($256\times 256$ versus $287$) when compared to the nonspatial component. The screen component further encodes important spatial information on the terrain and location of troops that has a significant impact on the policy network’s prediction.
@@
5.2.5 Adversarial Robustness of the Training Algorithm

### v2

@@ Abstract
keywords:
Adversarial Attacks, Deep Reinforcement Learning, Command and Control, Starcraft II, Adversarial Robustness
@@ 1 Introduction
Deep Reinforcement Learning (DRL) has been successfully used to train agents in several tactical and real-time strategy games such as StarCraftstarcraft and DoTAdota , which involve complex planning and decision-making. These agents have demonstrated proficiency in coming up with winning strategies comparable to that of experienced human players (AlphaStar vinyals , OpenAI Fiveopenai2019dota ) through techniques like self–play, imitation learning, etc. As a result, in recent years, there has been mounting interest in the military research community in applying these RL techniques to tasks such as operational planning and command and control (C2). Simultaneously, traditional game engines have been repurposed to facilitate automated learning (pySC2pysc2 , SMACsamvelyan2019starcraft ; smacgithub , pyDoTApydota2 ) and new ones developed for battlefield simulation narayanan2021first ; park2022deep ; soleyman2020multi ; zhang2020air ; BZCFHSWA2022 , creating what are effectively digital wargames. The driving force behind this research has been to improve and augment strategies used on the battlefields of the future, which are expected to be more complex and unconventional, possibly beyond the cognitive abilities of a human commander.
Recent works waytowich2022learning have had considerable success in winning simulated wargames using C2 agents that have been trained through reinforcement learning techniques and synthetic data. This has been possible partly due to the scalability of RL training which has proved to be a massive advantage for exploring and exploiting different strategies, even when faced with difficult or complicated scenarios and given only partial information about the environment. However, these evaluations are done in benign environments where information available to the C2 agent is assumed to be uncorrupted. Realistically, this is unlikely in situations on the battlefield where information may have inherent noise because of the mode of collection (from sensors or other sources) or may be tampered with by the enemy. In this work, we evaluate the robustness of such a trained agent when subject to potentially adversarial inputs in the context of C2.
@@
In order to do so, first, we use the StarCraft II Learning Environment (SCLEvinyals2017starcraft ) to model conflict between two teams, the BlueForce and the RedForce. The C2 agent directs the BlueForce to win the battle by eliminating the RedForce troops.
Next, we assume the an attacker present in the environment tampers with the observations collected from the battlefield before they are made available to the C2 agent. The added perturbations termed adversarial perturbations are constructed to be highly imperceptible to evade detection while maximally subverting the C2 agent’s policy to something deleterious (Figure 1). We then evaluate the resulting drop in performance of the agent on several metrics as well as analyze the deviation in the course of action from a military perspective.
@@
Our evaluations demonstrate the susceptibility of vanilla RL training algorithms to adversarial perturbations and the need for robust training mechanisms and sophisticated detection and prevention techniques especially for such critical scenarios.
@@ 2.1 RL for Command and Control (C2)
Mission success in military C2 requires disseminating control actions to designated forces based on real-time Intelligence, Surveillance, and Reconnaissance (ISR) information of the operational environment, such as kinetic and non-kinetic (e.g., weather, political, economic, cultural) variables, and terrain information. Before the battle begins, the commander and their specialized staffing officers must develop a detailed mission plan encapsulated as a Course Of Action (COA). The development of a COA requires a detailed analysis of the operational environment, predictions of the opposing force’s COA (strategy), and wargaming to identify a friendly force COA that is finely tuned to meet the mission requirements marr2001military . Typical military planning is solely based on the commander and staffing officers, and their allocated time before the battle dictates the number of possible COAs that can be considered. Additionally, each potential COA must be wargamed and fine-tuned against a small set of opposing force COAs to identify strengths and weaknesses. However, this can result in suboptimal (heuristic-based) COA s2000 .
To circumvent this limitation, future military planning is envisioned to incorporate an Artificial Intelligent (AI) commander’s assistant that can generate and recommend COAs to aid in the military planning process. Recent developments in deep RL for strategy games provides a promising direction to develop control policies for C2 vinyals . The algorithms developed allow an AI agent to learn the best control policy that optimizes a predefined reward function by playing millions of simulated games through the exploration of many environmental state and action pairs. To extend these formulations for military C2 requires modeling, simulating, and wargaming a large number of battles faster than real-time in a virtual environment that emulates realistic combat characteristics. Furthermore, the existing RL algorithms must be adapted to handle a large number of heterogeneous actors, doctrine based control strategies, complex state and action space, and uncertain information.
Previous work on RL for C2 found that the StarCraft II gaming engine, developed by Blizzard Entertainment starcraft , provides a simulation environment that can be militarized and used for prototyping an AI commander’s assistant goecks2023games ; waytowich2022learning . The following subsections provide details of the StarCraft II environment and the scenarios considered in this paper.
@@ 2.2 StarCraft II C2 Environment
StarCraft II is a multi-agent real-time strategy game developed by Blizzard Entertainment starcraft that consists of multiple players competing for influence and resources with the ultimate goal of defeating the other players. As a collaboration between Deepmind and Blizzard Entertainment, the StarCraft II Learning Environment (SC2LE) Machine Learning API was developed to allow researcher to study many difficult challenges associated with RL pysc2 . For example, controlling heterogeneous assets in a complex state and action space with uncertainties. Furthermore, the StarCraft II Editor allows developers to construct custom scenarios, making it possible to develop RL agents for C2.
Previous work extended this framework to militarize the SC2LE and develop a C2 simulation and experimentation capability that interfaces with deep RL algorithms via RLlib goecks2023games , an open-source industry-standard RL library liang2018rllib . The icons were re-skinned to portray standard military symbology, new StarCraft II maps were designed to emulate realistic combat characteristics, such as environment terrain and asset attributes (e.g., visibility, weapons, weapons ranges, and damage), and a custom wrapper was designed to train RL agents. For a detailed description of this framework, see goecks2023games . In our work, we use this framework to develop baseline RL agents for C2 and study the effects of adversarial attacks on the learned policies in the following two scenarios.
@@
TigerClaw. The TigerClaw melee map (Figure 3) is a high-level recreation of the TigerClaw combat scenario (Figure 2) developed using the StarCraft II map editor. The scenario was developed by Army subject-matter experts (SMEs) at the Captain’s Career Course, Fort Moore, Georgia. The BlueForce is an Armored Task Force (TF) which consists of combat armor with M1A2 Abrams, mechanized infantry with Bradley Fighting Vehicles (BFV), mortar, armored recon cavalry with BFV, and combat aviation. The RedForce is a Battalion Tactical Group (BTG) with attached artillery battery and consists of mechanized infantry with BMP, mobile artillery, armored recon cavalry, combat aviation, anti-armor with anti-tank guided missiles (ATGM), and combat infantry.
@@
National Training Center (NTC). The NTC map is a representation of the Ft. Irwin and Bicycle Lake area as seen in Figure 4. The Blue and RedForce units are the same as TigerClaw, but there is also a capability to randomize the starting forces. RedForce initially begins in the Bicycle Lake region, while BlueForce is set up in defense of Fort Irwin. Thus the RedForce will maneuver from the Bicycle Lake area and attempt to destroy the BlueForce in the Ft. Irwin area. The goal of this map was to investigate the impact of new terrain and to determine if the BlueForce would adopt a defensive strategy.
@@
Reinforcement Learning Algorithms
Asynchronous Advantage Actor-Critic (A3C) mnih2016asynchronous . A3C is an asynchronous version of the Advantage Actor-Critic (A2C) algorithm for RL that uses multiple agents to learn a policy and an estimate of the value function in parallel. At each timestep, the agent takes an action based on the current policy and receives a reward while transitioning to a new state. It then calculates the advantage function and updates its local copies of the actor-critic networks. This process takes place independently and asynchronously for each agent. A central parameter server that stores the global parameters of the networks is updated periodically by each agent and is used to initialize the local parameters of the agents’ local networks. The asynchronous nature of the updates leads to more efficient exploration of the environment and reduces the correlation between the updates leading to more stable and efficient learning. Typically, the policy and value functions are parameterized by a shared neural network with a softmax output for the policy and a linear output for the value function.
Proximal Policy Optimization (PPO) schulman2017proximal . PPO is a policy gradient method that aims to improve the performance and stability of trust-regionschulman2015trust methods by introducing a clipped surrogate objective function. This objective function effectively restricts policy change to a small range thus reducing the variability in training of the actor. During training, the PPO algorithm iteratively updates the actor as well as the action value function and state value function using the temporal-difference (TD) method. Further, this objective function enables PPO to guarantee monotonic improvements in the objective. This allows for faster convergence without strict constraints leading to more accurate and stable performance of the agent.
@@ 3.1 State and Action Space
The action space within the StarCraft II environment is large since it is a combination of the following three components: the number of units available to the commander, the number of possible actions that each unit can execute, and the $(x,y)$ pixel location within the size $256\times 256$ minimap where the action will be executed. To reduce the action space, the StarCraft II C2 environment first restricts the number of units by defining control groups, which lumps common units together to reduce the overall number of units needed to be controlled. In our custom scenarios, the control groups are defined as – 1 BlueForce: “AVIATION”, “MECH_INF”, “MORTAR”, “SCOUT” and “TANK2 RedForce: “ANTI_ARMOR”, “ARTILLERY”, “AVIATION”, “INFANTRY” and “MECH_INF
Within each control group, we restrict the possible actions to be either “NO_OP” or “ATTACK(x,y)”, where the function of the former is to essentially do nothing, while the latter moves the control group (i.e., all associated units) to a desired $(x,y)$ location and attacks any enemy within its firing range along the way. Additionally, the number of $(x,y)$ pixel locations is reduced by segmenting the minimap into nine disjoint quadrants with locations defined as “LEFT”, “CENTER”, and “RIGHT” for the $x$-axis, and “TOP”, “CENTER”, and “BOTTOM” for the $y$-axis, where the exact pixel location is the center of the quadrant.
Reward Structure
TigerClaw. The reward function for the TigerClaw map consists of -
@@
NTC. The NTC map shares the same attrition reward function as TigerClaw, but the terrain rewards are not included. That is,
@@ 3.2 RL Agent Description
The C2 agent in this paper uses the same policy network as our previous work waytowich2022learning . The network takes as inputs, three kinds of observations and the control group
@@ 4 Adversarial Attacks on RL agents
Prior works on the robustness of RL training have focused on evaluating the algorithms from a perspective of sensitivity to environment dynamics huang2017adversarial ; sun2020stealthy or the ability to train adaptive adversarial policies against them wu2021adversarial ; gleave2019adversarial . In this work, we focus on the former approach. A lot of previous research has shown that neural network predictions are highly sensitive to perturbations in their input space7958570 ; goodfellow2014explaining ; madry2018towards . As DRL approaches typically rely on parameterizing policies with neural networks, they suffer from the same vulnerabilities.
@@ 4.1 Adversarial Attacks on Image Classifiers
Fast Gradient Sign Method (FGSM) goodfellow2014explaining .
@@
where $y_{t}$ is the ground truth label of the sample $x$ and $\varepsilon$ is the perturbation budget that controls the amount of distortion to the original sample. In the case of FGSM, it bounds the $l_{\infty}$ norm of the perturbation added to the original image. Intuitively, the attack moves the sample $x$ in the direction of $\nabla_{x}L(f(w;x),y_{t})$ which maximizes the classifier loss $L$. In this work, we prefer FGSM to more powerful attacks such as PGDmadry2018towards and C&W7958570 because of its lower computational cost, which leads to a more efficient attack.
@@ 4.2 Inference Time Attacks on Policies
Attack Methodology.
We use FGSM as a basis for generating perturbations at inference time and use the modification presented by Huang et al.huang2017adversarial in order to target policies instead of classifiers. Unlike supervised learning, we do not know the “ground truth” action at any given timestep in RL and we assume that the action predicted by the policy network with the highest likelihood is optimal.
@@ 5.1 Experimental Setup
Our experiments were conducted on a cluster node with two AMD Epyc 7763 “Milan” CPUs 128 cores with 256GBs of main memory. We used the SC2 framework described in Section 2.2 to train our RL agents. Each training run was performed using 90 workers, each occupying a single core (90 parallel instances of the game) and consumed about 170GBs of memory. The PPO and A3C agents were trained for 5K/25K iterations respectively corresponding to around 40M/100M timesteps or 134K/376K episodes respectively. The rollouts were visualized using a custom pygame interface.
@@ 5.2 Evaluation of Inference Time Attacks on Agent Reward
Vulnerability to Adversarial Perturbations.
@@
Analyzing Agent Behavior under Attack.
@@
To better understand this shift in behavior, we compare actions taken by the agent in a benign environment to those taken by the same agent when an adversary is present. That is, at each timestep we compute the actions that would be taken by the agent if the observations were maliciously modified and compare these “subverted actions” to the actions that are actually taken by the agent at that timestep. We do so by perturbing the observations using our attack (with $\varepsilon=0.1$) and using the agent to predict the next action. The normalized action frequencies are plotted in Figure 10 for a PPO/TigerClaw agent over 5 episodes. A clear difference is observed in the actions taken in the two cases. Actions $(0,2,1),(1,1,1),(1,2,1)$ are sampled frequently when perturbed observations are presented to the agent. These actions misdirect the BlueForce as they target areas on the map that do not contain any RedForce troops.
@@
Strength & Reliability of the Attack.
@@
Attacking the A3C/TigerClaw agent. As can be seen in Figure 7(b), the attack appears to fail on the C2 agent trained using A3C on the TigerClaw map. Even for large perturbations ($\varepsilon>0.1$), the attack fails to degrade the agent’s reward as extensively as the other cases. Interestingly, we also note that even in a benign environment, the A3C/TigerClaw agent achieves a significantly lower reward than the PPO/TigerClaw agent.
@@
Component-wise Impact of Input on Attack Results.
@@
Figures 17(a), 17(b), and 17(d) indicate that for smaller perturbations levels ($\epsilon<0.5$), modifications to the screen component are primarily responsible for a successful attack. Further, we only see a minor degradation in the reward for small nonspatial-only perturbations in both scenarios. On the other hand, the drop in mean reward because of screen-only perturbations is similar to that in the baseline attack. Intuitively, this skewed importance makes sense as the screen component is a much larger input component ($256\times 256$ versus $287$) when compared to the nonspatial component. The screen component further encodes important spatial information on the terrain and location of troops that has a significant impact on the policy network’s prediction.
@@
Adversarial Robustness of the Training Algorithm.